40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5749     Entry time: Fri Oct 28 01:13:17 2011
Author: kiwamu 
Type: Update 
Category: SUS 
Subject: ITMX oplev : iris fully opened 

I found that the sum of the ITMX oplev signals had gone down to zero yesterday.

I checked the ITMX table and found two iris on the He-Ne laser path were blocking the beam on their apertures.

I guess this is because we were working around there for installation of POP/POX and may have touched some of the oplev optics.

Then I fully opened the apertures of those two iris and the sum went back to nominal of 600 counts.

ELOG V3.1.3-